### v1

# Probabilistic Variational Contrastive Learning

Abstract
Deterministic embeddings learned by contrastive learning (CL) methods such as SimCLR and SupCon achieve state-of-the-art performance but lack a principled mechanism for uncertainty quantification. We propose Variational Contrastive Learning (VCL), a decoder-free framework that maximizes the evidence lower bound (ELBO) by interpreting the InfoNCE loss as a surrogate reconstruction term and adding a KL divergence regularizer to a uniform prior on the unit hypersphere. We model the approximate posterior $q_{\theta}({\boldsymbol{z}}|{\boldsymbol{x}})$ as a projected normal distribution, enabling the sampling of probabilistic embeddings. Our two instantiations—VSimCLR and VSupCon—replace deterministic embeddings with samples from $q_{\theta}({\boldsymbol{z}}|{\boldsymbol{x}})$ and incorporate a normalized KL term into the loss. Experiments on multiple benchmarks demonstrate that VCL mitigates dimensional collapse, enhances mutual information with class labels, and matches or outperforms deterministic baselines in classification accuracy, all the while providing meaningful uncertainty estimates through the posterior model. VCL thus equips contrastive learning with a probabilistic foundation, serving as a new basis for contrastive approaches.
1 Introduction
Deep representation learning seeks to map each input ${\boldsymbol{x}}$ into a compact embedding ${\boldsymbol{z}}$ that preserves semantic similarity and facilitates downstream tasks such as classification or retrieval bengio2013representation . Contrastive learning methods, including SimCLR chen2020simple and SupCon khosla2020supervised , have advanced the state of the art by pulling together positive pairs and pushing apart negatives in the embedding space. However, these approaches rely on deterministic point estimates for each sample, which do not express uncertainty or capture multiple plausible representations.
To address this limitation, we introduce a probabilistic Variational Contrastive Learning (VCL) approach, which extends deterministic embeddings to probabilistic embeddings by maximizing the evidence lower bound (ELBO) within the contrastive learning framework. Unlike variational autoencoders (VAEs) kingma2019introduction , which employ a decoder to reconstruct inputs from latent variables, VCL omits explicit decoders. Instead, we show that the InfoNCE loss can serve as a surrogate for the ELBO reconstruction term, yielding a principled probabilistic formulation of contrastive learning. Our VCL framework offers several new perspectives on learned embeddings:
Understanding embeddings through distributions.
VCL maps each input ${\boldsymbol{x}}$ to an approximate posterior distribution $q_{\phi}({\boldsymbol{z}}|{\boldsymbol{x}})$, yielding a mean vector that serves as the embedding and a variance that quantifies uncertainty. This probabilistic representation not only captures richer information about each sample but also enables uncertainty-aware downstream decisions.
A probabilistic ELBO viewpoint beyond mutual information.
Minimizing the InfoNCE loss maximizes a lower bound on mutual information, consistent with the InfoMax principle linsker1988self . However, mutual information alone may not directly correlate with downstream task performance Tschannen2020On , motivating geometric analyses based on alignment and uniformity wang2020understanding . In contrast, our probabilistic formulation interprets contrastive learning through the ELBO: maximizing the ELBO with a fixed decoder encourages the learned posterior $q_{\phi}({\boldsymbol{z}}|{\boldsymbol{x}})$ to match the true posterior $p({\boldsymbol{z}}|{\boldsymbol{x}})$, providing a principled objective for representation learning.
Controllable embedding distributions.
Standard contrastive learning imposes no explicit prior on embeddings, so their distribution emerges implicitly from the data and model architecture. VCL incorporates a prior $p({\boldsymbol{z}})$ into the objective, allowing one to specify and control the marginal embedding distribution $p({\boldsymbol{z}})$. For example, choosing a uniform prior on the hypersphere often improves empirical performance wang2020understanding that makes negative samples more uniformly dispersed around each anchor, and domain-specific priors can encode known structure in the data.
Mitigating collapse phenomena.
Self-supervised contrastive learning (e.g., SimCLR) can suffer from dimensional collapse, where embeddings occupy only a low-dimensional subspace jing2021understanding . A spherical uniform prior mitigates this by encouraging isotropic use of all embedding dimensions.
This Variational Contrastive Learning framework thus provides uncertainty-aware embeddings, a new basis of CL with theoretical insights via the ELBO, and practical solutions to common collapse problems in contrastive learning. Our contributions are summarized as follows:
  • We introduce Variational Contrastive Learning (VCL), a decoder-free ELBO maximization framework that reinterprets the InfoNCE loss as a surrogate reconstruction term and incorporates a KL divergence regularizer to a uniform prior on the unit hypersphere.
  • We propose a distributional embedding model using a projected normal posterior $q_{\theta}({\boldsymbol{z}}|{\boldsymbol{x}})$ that enables sampling, uncertainty quantification, and efficient KL computation on the hypersphere.
  • We derive a theoretical connection between the optimal InfoNCE critic and the ELBO, showing that minimizing InfoNCE asymptotically maximizes the ELBO reconstruction term (Proposition 3.1) and providing a generalization bound (Theorem 3.2) of KL regularization.
  • We demonstrate that VCL mitigates both dimensional collapse in self-supervised contrastive learning via the KL regularizer, while preserving embedding structure.
  • Through experiments on CIFAR-10, CIFAR-100, STL-10, Tiny-ImageNet, and Caltech-256, we show that VCL methods (VSimCLR and VSupCon) preserve or improve mutual information with labels, match or exceed classification accuracy of deterministic baselines, and provide meaningful implication of distributional embeddings.
2 Preliminaries
Let $\mathcal{D}=\{({\boldsymbol{x}}_{i},{\boldsymbol{y}}_{i})\}_{i=1}^{N}$ be a dataset of input ${\boldsymbol{x}}\in\mathcal{X}$ and label pairs drawn i.i.d. from the joint distribution $p({\boldsymbol{x}},{\boldsymbol{y}})$. An encoder $f_{\theta}\colon\mathcal{X}\to\mathbb{R}^{d}$, parameterized by $\theta$, maps each input ${\boldsymbol{x}}$ to a $d$-dimensional vector, which we then normalize to unit length: ${\boldsymbol{z}}=\frac{f_{\theta}({\boldsymbol{x}})}{\|f_{\theta}({\boldsymbol{x}})\|_{2}}$. Throughout this section, we define the temperature–scaled cosine similarity between embeddings ${\boldsymbol{z}}_{i}$ and ${\boldsymbol{z}}_{j}$ as
$s({\boldsymbol{z}}_{i},{\boldsymbol{z}}_{j})=\frac{{\boldsymbol{z}}_{i}^{\top}{\boldsymbol{z}}_{j}}{\tau},$ (1)
where $\tau>0$ is the temperature hyperparameter. For any two probability distributions $q$ and $p$, we denote the Kullback–Leibler (KL) divergence by $D(q\,\|\,p)\;=\;\mathbb{E}_{{\boldsymbol{z}}\sim q}\!\left[\log\frac{q({\boldsymbol{z}})}{p({\boldsymbol{z}})}\right]$.
2.1 Self-Supervised Contrastive Learning
Self-supervised contrastive learning (SSCL) learns representations from unlabeled data by pulling together embeddings of semantically related views (positives) and pushing apart those of unrelated views (negatives). For an anchor ${\boldsymbol{x}}$, let ${\boldsymbol{x}}^{\prime}_{i}$ denote a positive view sampled from $p({\boldsymbol{x}}^{\prime}_{i}\mid{\boldsymbol{x}})$, and let $\{{\boldsymbol{x}}^{\prime}_{j}\}_{j\neq i}$ be $N-1$ negative views drawn i.i.d. from the marginal $p({\boldsymbol{x}}^{\prime})$. The InfoNCE loss oord2018representation for anchor ${\boldsymbol{x}}$ is then
$I_{\mathrm{NCE}}({\boldsymbol{x}};{\boldsymbol{x}}^{\prime})$ $=-\mathbb{E}_{\begin{subarray}{c}{\boldsymbol{x}}\sim p({\boldsymbol{x}})\\ {\boldsymbol{x}}^{\prime}_{i}\sim p({\boldsymbol{x}}^{\prime}_{i}\mid{\boldsymbol{x}})\\ \{{\boldsymbol{x}}^{\prime}_{j}\}_{j\neq i}\sim p({\boldsymbol{x}}^{\prime})\end{subarray}}\left[\log\frac{\exp\bigl{(}s({\boldsymbol{z}},{\boldsymbol{z}}^{\prime}_{i})\bigr{)}}{\sum_{j=1}^{N}\exp\bigl{(}s({\boldsymbol{z}},{\boldsymbol{z}}^{\prime}_{j})\bigr{)}}\right],$ (2)
where ${\boldsymbol{z}}=f_{\theta}({\boldsymbol{x}})/\|f_{\theta}({\boldsymbol{x}})\|_{2}$ and $s(\cdot,\cdot)$ is the temperature-scaled cosine similarity.
In practice, following SimCLR chen2020simple , we generate positives by applying two random augmentations $t^{\prime},t^{\prime\prime}\sim\mathcal{T}$ to each sample ${\boldsymbol{x}}_{i}$, yielding $({\boldsymbol{x}}^{\prime}_{i},{\boldsymbol{x}}^{\prime\prime}_{i})=(t^{\prime}({\boldsymbol{x}}_{i}),\,t^{\prime\prime}({\boldsymbol{x}}_{i}))$. [^1] Although we adopt the SimCLR augmentation scheme, our method applies to any contrastive framework. All other $22N-$ 2 augmented samples in the mini-batch serve as negatives. Let $\mathcal{B}$ be the set of all $2N$ embeddings in the batch; then InfoNCE can be computed as
$I_{\mathrm{NCE}}$ $=-\frac{1}{2N}\sum_{{\boldsymbol{z}}\in\mathcal{B}}\log\frac{\exp\bigl{(}s({\boldsymbol{z}},{\boldsymbol{z}}_{p})\bigr{)}}{\sum_{{\boldsymbol{z}}_{n}\in\mathcal{B}\setminus\{{\boldsymbol{z}}\}}\exp\bigl{(}s({\boldsymbol{z}},{\boldsymbol{z}}_{n})\bigr{)}},$ (3)
where ${\boldsymbol{z}}_{p}$ denotes the positive embedding corresponding to ${\boldsymbol{z}}$. Since InfoNCE lower-bounds the mutual information $I({\boldsymbol{x}};{\boldsymbol{x}}^{\prime})$ via
$I({\boldsymbol{x}};{\boldsymbol{x}}^{\prime})\;\geq\;\log N\;-\;I_{\mathrm{NCE}}({\boldsymbol{x}};{\boldsymbol{x}}^{\prime}),$ (4)
we can see that minimizing $I_{\mathrm{NCE}}$ encourages encoders to preserve the semantic information of ${\boldsymbol{x}}$ poole2019variational .
2.2 Supervised Contrastive Learning
Khosla et al. khosla2020supervised extend the InfoNCE loss from the self-supervised setting to a supervised context, calling the resulting method Supervised Contrastive Learning (SupCon). When class labels $y_{i}\in\{1,\dots,C\}$ are available, all samples sharing the same label can serve as positives.
Given a mini-batch $\{({\boldsymbol{x}}_{i},y_{i})\}_{i=1}^{B}$, define for each anchor index $a$
$\mathcal{A}(a)=\{1,2,\dots,B\}\setminus\{a\},\text{ and }~{}\mathcal{P}(a)=\{\,p\in\mathcal{A}(a):y_{p}=y_{a}\},$
so that $\mathcal{P}(a)$ contains the indices of all positives for anchor $a$. The SupCon loss for anchor ${\boldsymbol{x}}_{a}$ is then
$I_{\mathrm{SUP}}({\boldsymbol{x}}_{a})$ $=-\frac{1}{|\mathcal{P}(a)|}\sum_{p\in\mathcal{P}(a)}\log\frac{\exp\bigl{(}s({\boldsymbol{z}}_{a},{\boldsymbol{z}}_{p})\bigr{)}}{\displaystyle\sum_{j\in\mathcal{A}(a)}\exp\bigl{(}s({\boldsymbol{z}}_{a},{\boldsymbol{z}}_{j})\bigr{)}}.$ (5)
Averaging over all anchors in the batch yields the full objective:
$\mathcal{L}^{\mathrm{sup}}$ $=\frac{1}{B}\sum_{a=1}^{B}I_{\mathrm{SUP}}({\boldsymbol{x}}_{a}).$ (6)
2.3 Variational Inference and the Evidence Lower Bound (ELBO)
In variational inference blei2017variational ; kingma2019introduction , we treat the data distribution $p({\boldsymbol{x}})$ as the marginal of a joint distribution over observed data ${\boldsymbol{x}}$ and latent variables ${\boldsymbol{z}}$, i.e., $p({\boldsymbol{x}})=\int p({\boldsymbol{x}}|{\boldsymbol{z}})\,p({\boldsymbol{z}})\mathrm{d}{\boldsymbol{z}}.$ The latent variable ${\boldsymbol{z}}$ captures meaningful structure in ${\boldsymbol{x}}$, serving both as a hidden cause and as a compressed representation for downstream tasks. In representation learning, we interpret ${\boldsymbol{z}}$ as the embedding of ${\boldsymbol{x}}$.
The log-evidence can be written with respect to any approximate posterior $q_{\phi}({\boldsymbol{z}}|{\boldsymbol{x}})$ as
$\log p({\boldsymbol{x}})$ $=\log\mathbb{E}_{q_{\phi}({\boldsymbol{z}}\mid{\boldsymbol{x}})}\!\Bigl{[}\tfrac{p({\boldsymbol{x}},{\boldsymbol{z}})}{q_{\phi}({\boldsymbol{z}}\mid{\boldsymbol{x}})}\Bigr{]}.$ (7)
Rather than optimizing (7) directly, variational methods maximize the evidence lower bound (ELBO) obtained as a result of applying Jensen’s inequality:
$\log p({\boldsymbol{x}})$ $\geq\mathbb{E}_{q_{\phi}({\boldsymbol{z}}\mid{\boldsymbol{x}})}\!\bigl{[}\log p({\boldsymbol{x}}|{\boldsymbol{z}})\bigr{]}-D\bigl{(}q_{\phi}({\boldsymbol{z}}|{\boldsymbol{x}})\,\|\,p({\boldsymbol{z}})\bigr{)}\;=\;\mathcal{L}^{\rm ELBO}(\phi),$ (8)
where $p({\boldsymbol{z}})$ is a fixed prior (commonly $\mathcal{N}(\mathbf{0},I_{d})$). The ELBO decomposes into a reconstruction term $\mathbb{E}_{q}[\log p({\boldsymbol{x}}|{\boldsymbol{z}})]$ and a regularizer $D(q_{\phi}({\boldsymbol{z}}|{\boldsymbol{x}})\,\|\,p({\boldsymbol{z}}))$. Maximizing $\mathcal{L}^{\rm ELBO}$ thus balances (i) accurate reconstruction, (ii) posterior-to-prior regularization, and (iii) posterior accuracy. By
$\log p({\boldsymbol{x}})=\mathcal{L}^{\rm ELBO}(\phi)+D\bigl{(}q_{\phi}({\boldsymbol{z}}|{\boldsymbol{x}})\,\|\,p({\boldsymbol{z}}|{\boldsymbol{x}})\bigr{)},$ (9)
for fixed $\log p({\boldsymbol{x}})$, maximizing the ELBO minimizes the KL divergence between the approximate and true posteriors blei2017variational .
The ELBO provides a tractable surrogate for marginal likelihood that can be optimized by standard gradient methods. It will serve as the theoretical backbone of our Variational Contrastive Learning framework, offering both a probabilistic interpretation and explicit control over latent uncertainty.
Relation to contrastive objectives.
Although the ELBO stems from latent-variable modeling, its two components align naturally with contrastive objectives: the KL divergence term enforces uniformity in the embedding space, while the reconstruction term promotes alignment between embeddings and observations. In Section 3, we leverage this connection by adopting distributional embeddings in the contrastive framework and incorporating a KL-based regularizer on the posterior.
3 Variational Contrastive Learning (VCL)
Unlike existing variational contrastive learning methods—which primarily focus on generative models with explicit decoders chen2025multi ; wang2024vcl —our approach performs decoder-free ELBO maximization, making VCL a truly contrastive framework.
3.1 Decoder-Free ELBO Maximization
Here we describe how to optimize two terms in ELBO (8) within a purely contrastive setup.
Reconstruction term.
The reconstruction term $\mathbb{E}_{q_{\theta}({\boldsymbol{z}}|{\boldsymbol{x}})}\bigl{[}\log p({\boldsymbol{x}}|{\boldsymbol{z}})\bigr{]}$ requires the true conditional $p({\boldsymbol{x}}|{\boldsymbol{z}})$, which is generally intractable. Instead, we approximate it via the embedding conditional
$p({\boldsymbol{z}}^{\prime}|{\boldsymbol{z}})=\frac{p({\boldsymbol{z}},{\boldsymbol{z}}^{\prime})}{\int p({\boldsymbol{z}},{\boldsymbol{z}}^{\prime})\,\mathrm{d}{\boldsymbol{z}}^{\prime}},$ (10)
where ${\boldsymbol{z}}^{\prime}\sim q_{\theta}(\cdot\mid{\boldsymbol{x}})$ captures semantics of ${\boldsymbol{x}}$. Thus,
$\mathbb{E}_{q_{\theta}({\boldsymbol{z}}|{\boldsymbol{x}})}\bigl{[}\log p({\boldsymbol{x}}|{\boldsymbol{z}})\bigr{]}$ $\approx\mathbb{E}_{q_{\theta}({\boldsymbol{z}},{\boldsymbol{z}}^{\prime}|{\boldsymbol{x}})}\bigl{[}\log p({\boldsymbol{z}}^{\prime}|{\boldsymbol{z}})\bigr{]}$
$=\mathbb{E}\Bigl{[}\log\tfrac{p({\boldsymbol{z}},{\boldsymbol{z}}^{\prime})}{\int p({\boldsymbol{z}},{\boldsymbol{z}}^{\prime})\,\mathrm{d}{\boldsymbol{z}}^{\prime}}\Bigr{]}\approx\mathbb{E}\Bigl{[}\log\tfrac{e^{\psi({\boldsymbol{z}},{\boldsymbol{z}}^{\prime})}}{\sum_{j}e^{\psi({\boldsymbol{z}},{\boldsymbol{z}}^{\prime}_{j})}}\Bigr{]},$ (11)
where we approximate $p({\boldsymbol{z}},{\boldsymbol{z}}^{\prime})\!\approx\!e^{\psi({\boldsymbol{z}},{\boldsymbol{z}}^{\prime})}$ via a critic $\psi$ and use a Monte Carlo estimate. Details on parameterizing $p({\boldsymbol{z}}^{\prime}\mid{\boldsymbol{z}})$ appear in Section 3.2.
Noting that the right-hand side of (3.1) is (up to sign) the InfoNCE surrogate, setting $\psi(\cdot,\cdot)=s(\cdot,\cdot)$ in (3.1) where $s(\cdot,\cdot)$ is defined in (1) yields
$\mathbb{E}_{q_{\theta}({\boldsymbol{z}}|{\boldsymbol{x}})}\bigl{[}\log p({\boldsymbol{x}}|{\boldsymbol{z}})\bigr{]}\approx-\,I_{\mathrm{NCE}}({\boldsymbol{x}};{\boldsymbol{x}}^{\prime}).$ (12)
Hence, minimizing the InfoNCE loss maximizes the ELBO’s reconstruction term without any explicit decoder.
In contrast to VAE embeddings—which often rely on pixel-level reconstruction through expressive decoder song2024scale —VCL preserves semantics via contrastive objectives. The next proposition provides a theoretical connection between InfoNCE and the reconstruction term.
Proposition 3.1.
Assume the critic $\psi$ in InfoNCE is optimal. As the number of negatives $N\to\infty$,
$-\,I_{\mathrm{NCE}}({\boldsymbol{x}};{\boldsymbol{x}}^{\prime})+\log N\;\longrightarrow\;\mathbb{E}\bigl{[}\log p({\boldsymbol{z}}^{\prime}|{\boldsymbol{z}})\bigr{]}-D\bigl{(}q_{\theta}({\boldsymbol{z}}^{\prime}|{\boldsymbol{x}})\,\|\,p({\boldsymbol{z}}^{\prime})\bigr{)}-H\bigl{(}q_{\theta}({\boldsymbol{z}}^{\prime}|{\boldsymbol{x}})\bigr{)},$ (13)
where the expectation is over $q_{\theta}({\boldsymbol{z}}|{\boldsymbol{x}})q_{\theta}({\boldsymbol{z}}^{\prime}|{\boldsymbol{x}})$, and $H(\cdot)$ denotes entropy.
Proof.
The proof of Proposition 3.1 is in Appendix B.1. ∎
Regularization.
Maximizing the ELBO requires choosing a prior $p({\boldsymbol{z}})$ and an approximate posterior $q_{\theta}({\boldsymbol{z}}\mid{\boldsymbol{x}})$. Although both are often taken as Gaussian distributions kingma2019introduction , this choice conflicts with the geometry of contrastive embeddings, which often lie on the unit hypersphere due to the normalization wang2020understanding . Instead, we adopt non-Gaussian priors and posteriors—one key distinction from standard VAE approaches.
Motivated by the uniformity property wang2020understanding on the unit sphere $\mathcal{S}^{d-1}=\{{\boldsymbol{z}}\in\mathbb{R}^{d}:\|{\boldsymbol{z}}\|_{2}=1\}$, we set the prior $p({\boldsymbol{z}})$ to be the uniform distribution over $\mathcal{S}^{d-1}$. For the approximate posterior, we use the projected normal distribution hernandez2017general , which admits efficient KL-divergence computation while enforcing ${\boldsymbol{z}}\in\mathcal{S}^{d-1}$. A random variable ${\boldsymbol{z}}\sim\mathcal{P}\mathcal{N}(\mu,K)$ is obtained by sampling
${\boldsymbol{z}}=\frac{{\bf u}}{\|{\bf u}\|_{2}}\quad\text{with}\quad{\bf u}\sim\mathcal{N}(\mu,K).$ (14)
[…]
With $q_{\theta}({\boldsymbol{z}}|{\boldsymbol{x}})=\mathcal{P}\mathcal{N}(\mu,K)$, the regularization term becomes
$D\bigl{(}q_{\theta}({\boldsymbol{z}}|{\boldsymbol{x}})\,\|\,p({\boldsymbol{z}})\bigr{)}=D\bigl{(}\mathcal{P}\mathcal{N}(\mu,K)\,\|\,\mathrm{Unif}(\mathcal{S}^{d-1})\bigr{)}.$ (15)
Since a closed-form KL divergence between projected normals and the uniform sphere is intractable, we instead minimize the Gaussian KL as an upper bound—by the data processing inequality polyanskiy2025information :
$D\bigl{(}\mathcal{N}(\mu,K)\,\|\,\mathcal{N}(0,I_{d})\bigr{)}\;\geq\;D\bigl{(}\mathcal{P}\mathcal{N}(\mu,K)\,\|\,\mathrm{Unif}(\mathcal{S}^{d-1})\bigr{)}.$ (16)
For $K=\mathrm{diag}(\sigma_{1}^{2},\dots,\sigma_{d}^{2})$, the Gaussian KL admits the closed form
$D(\mu,K)=\frac{1}{2}\sum_{i=1}^{d}\bigl{(}\sigma_{i}^{2}+\mu_{i}^{2}-1-\log\sigma_{i}^{2}\bigr{)}.$ (17)
The KL divergence term $D(\mu,K)$ grows linearly with the embedding dimension $d$, which can destabilize training when $d$ is large. To address this, we normalize the KL term by $d$, i.e., $\widetilde{D}(\mu,K)\;=\;\frac{1}{d}\,D(\mu,K),$ so that its magnitude remains comparable to the InfoNCE loss.
Final objective for maximizing ELBO.
By combining (12) and (17), we obtain the following (approximate) lower bound on the ELBO:
$\mathcal{L}^{\rm ELBO}(\theta)\;\geq\;-\,I_{\rm NCE}({\boldsymbol{x}};{\boldsymbol{x}}^{\prime})\;-\,D\bigl{(}\mu_{{\boldsymbol{x}}},\,K_{{\boldsymbol{x}}}\bigr{)},$ (18)
where $\mu_{{\boldsymbol{x}}}$ and $K_{{\boldsymbol{x}}}=\mathrm{diag}(\sigma_{{\boldsymbol{x}},1},\ldots,\sigma_{{\boldsymbol{x}},d})$ are the parameters of $q_{\theta}(\mathbf{z}\mid{\boldsymbol{x}})$. Because this bound is asymmetric in $({\boldsymbol{x}},{\boldsymbol{x}}^{\prime})$, we symmetrize it to define our final VCL objective:
$\mathcal{L}^{\rm VCL}=\frac{1}{2}\Bigl{(}I_{\rm NCE}({\boldsymbol{x}};{\boldsymbol{x}}^{\prime})+I_{\rm NCE}({\boldsymbol{x}}^{\prime};{\boldsymbol{x}})+D(\mu_{{\boldsymbol{x}}},K_{{\boldsymbol{x}}})+D(\mu_{{\boldsymbol{x}}^{\prime}},K_{{\boldsymbol{x}}^{\prime}})\Bigr{)}.$ (19)
Minimizing $\mathcal{L}^{\rm VCL}$ therefore maximizes the ELBO. Next, we introduce Variational SimCLR (VSimCLR), which is specifically designed to optimize this objective efficiently.
3.2 Variational SimCLR (VSimCLR)
We propose Variational SimCLR (VSimCLR), whose architecture is illustrated in Figure LABEL:subfig:VSimCLR. VSimCLR minimizes $\mathcal{L}^{\rm VCL}$ in (19), thereby implicitly maximizing the ELBO and bringing the approximate posterior closer to the true posterior by (9). Compared to SimCLR, VSimCLR differs in three key aspects: (i) the encoder outputs the parameters of a variational posterior rather than deterministic embeddings; (ii) embeddings are sampled from this posterior; and (iii) a KL divergence term between the approximate posterior and the prior is included in the loss.
Specifically, during training, each input ${\boldsymbol{x}}$ is first augmented twice to obtain ${\boldsymbol{x}}^{\prime}$ and ${\boldsymbol{x}}^{\prime\prime}$, as in SimCLR. The encoder then maps ${\boldsymbol{x}}^{\prime}$ and ${\boldsymbol{x}}^{\prime\prime}$ to posterior parameters $(\boldsymbol{\mu}^{\prime},\boldsymbol{\sigma}^{\prime})$ and $(\boldsymbol{\mu}^{\prime\prime},\boldsymbol{\sigma}^{\prime\prime})$, respectively. We then sample
${\boldsymbol{z}}^{\prime}=\boldsymbol{\mu}^{\prime}+\operatorname{diag}(\boldsymbol{\sigma}^{\prime})\;\boldsymbol{\epsilon}_{1},\quad\text{and}\quad{\boldsymbol{z}}^{\prime\prime}=\boldsymbol{\mu}^{\prime\prime}+\operatorname{diag}(\boldsymbol{\sigma}^{\prime\prime})\;\boldsymbol{\epsilon}_{2},$ (20)
where $\boldsymbol{\epsilon}_{1},\boldsymbol{\epsilon}_{2}\overset{\rm i.i.d}{\sim}\mathcal{N}(\bm{0},I_{d})$. After normalizing $\mathbf{z}^{\prime}$ and $\mathbf{z}^{\prime\prime}$ to unit length, we compute the InfoNCE loss over the normalized embeddings in the batch and add the KL divergence
$\frac{1}{d}D\bigl{(}\mathcal{N}(\boldsymbol{\mu},\operatorname{diag}(\boldsymbol{\sigma}^{2}))\,\|\,\mathcal{N}(\bm{0},I_{d})\bigr{)}$ (21)
for each sample. Minimizing this combined objective effectively minimizes $\mathcal{L}^{\rm VCL}$ in (19) and thus maximizes the ELBO. Figure LABEL:fig:VsimCLR highlights these differences: VSimCLR replaces deterministic embeddings with the projected-normal posterior $\mathcal{P}\mathcal{N}(\boldsymbol{\mu},\operatorname{diag}(\boldsymbol{\sigma}^{2}))$ and regularizes it via KL divergence to the standard normal.
3.3 Variational SupCon (VSupCon)
[…]
${\boldsymbol{z}}^{\prime}\sim\mathcal{P}\mathcal{N}(\hbox{\boldmath$\mu$}^{\prime},K^{\prime}),\qquad{\boldsymbol{z}}^{\prime\prime}\sim\mathcal{P}\mathcal{N}(\hbox{\boldmath$\mu$}^{\prime\prime},K^{\prime\prime}).$ (22)
Then the VSupCon objective is the symmetrized supervised loss plus the averaged, normalized KL penalties:
$\mathcal{L}^{\mathrm{VSup}}=\frac{1}{2}\Bigl{(}\mathcal{L}^{\rm sup}({\boldsymbol{z}}^{\prime},{\boldsymbol{z}}^{\prime\prime})+\mathcal{L}^{\rm sup}({\boldsymbol{z}}^{\prime\prime},{\boldsymbol{z}}^{\prime})\Bigr{)}\;+\;\frac{1}{2d}\Bigl{(}D(\hbox{\boldmath$\mu$}^{\prime},K^{\prime})+D(\hbox{\boldmath$\mu$}^{\prime\prime},K^{\prime\prime})\Bigr{)}.$ (23)
Minimizing $\mathcal{L}^{\mathrm{VSup}}$ therefore aligns same-class embeddings and regularizes their posterior distributions toward the uniform prior on the sphere.
3.4 Generalization analysis for KL divergence
Generalization bounds quantify how a loss function performs on unseen data. While recent work has extensively studied the InfoNCE loss saunshi2019theoretical ; lei2023generalization ; hieu2025generalization , the KL regularizer in VCL has not yet received comparable theoretical treatment. To address this gap, we derive a generalization bound for the KL term under the deep neural network encoder function class $\mathcal{F}_{\bm{\Xi}}$, parameterized by a bounded weight set $\bm{\Xi}$ and equipped with Lipschitz activation functions, as considered in the recent work hieu2025generalization . We introduce an informal version of the generalization bound below; the formal version is provided in Theorem B.1 in Appendix B.2.
Theorem 3.2 (Informal).
Let $\{\bm{x}_{i}\}_{i=1}^{N}\overset{\mathrm{i.i.d}}{\sim}p({\boldsymbol{x}})$, and let $D_{\mathrm{KL}}(f_{\bm{\Theta}};\bm{x})$ denote the KL regularizer applied to the output of the encoder $f_{\bm{\Theta}}\in\mathcal{F}_{\bm{\Xi}}$ given input $\bm{x}$. Then, with high probability, it holds:
$\sup_{f_{\bm{\Theta}}\in\mathcal{F}_{\bm{\Xi}}}\left[\mathbb{E}_{\bm{x}\sim p({\boldsymbol{x}})}\,D_{\rm KL}(f_{\bm{\Theta}};\bm{x})-\frac{1}{N}\sum_{i=1}^{N}D_{\rm KL}(f_{\bm{\Theta}};\bm{x}_{i})\right]\leq\tilde{\mathcal{O}}(1/\sqrt{N}).$ (24)
Theorem 3.2 shows that the generalization gap of the KL term decays as $\tilde{\mathcal{O}}(1/\sqrt{N})$. In contrast, as shown in (hieu2025generalization, , Theorem 1), the gap of InfoNCE scales as $\tilde{\mathcal{O}}(1)$ and does not improve with the number of negative samples $N$. Moreover, their analysis assumes that $\{\bm{x}_{i}\}_{i=1}^{N}$ are drawn from class-conditional distributions, which is a strictly stronger assumption than our i.i.d. assumption from the marginal distribution. Therefore, the KL regularizer does not degrade the generalization guarantees of InfoNCE, while also providing principled uncertainty quantification.
4 Experiments
We evaluate VCL with SimCLR and SupCon across five aspects: (i) embedding visualization, (ii) dimensional collapse, (iii) mutual information between embeddings and labels, (iv) classification accuracy, and (v) implications of distributional embeddings. Experiments are conducted on CIFAR-10 krizhevsky2009learning , CIFAR-100 krizhevsky2009learning , STL-10 coates2011analysis , Tiny-ImageNet le2015tiny , and Caltech-256 griffin2007caltech . We use ResNet-18 he2016deep as the encoder, set the embedding dimension to $d=128$, and employ a linear classifier for downstream evaluations. Additional implementation and training details are provided in Appendix C.
4.1 Embedding Visualization
Figure LABEL:fig:emb_visualization presents t-SNE van2008visualizing and UMAP mcinnes2018umap projections of the embeddings learned by SimCLR and VSimCLR on the CIFAR-10 test set. Although VSimCLR incorporates an additional KL-regularizer, it preserves the characteristic cluster structure induced by contrastive learning. This confirms that our distributional embeddings retain the semantic information learned by contrastive methods, rather than prioritizing reconstruction quality of decoders in VAE-based approaches.
4.2 Dimensional Collapse
Figure 2: Singular-value spectrum of the embedding covariance on CIFAR-10 and CIFAR-100. VSimCLR mitigates dimensional collapse.
Contrastive learning methods such as SimCLR often suffer from dimensional collapse, where embeddings concentrate in a low-dimensional subspace, underutilizing the full capacity of the representation space jing2021understanding . To quantify this effect, let $\{{\boldsymbol{z}}_{i}\}_{i=1}^{N}$ be the test-set embeddings and their covariance matrix $C=\frac{1}{N}\sum_{i=1}^{N}({\boldsymbol{z}}_{i}-\bar{{\boldsymbol{z}}})({\boldsymbol{z}}_{i}-\bar{{\boldsymbol{z}}})^{\top},$ with $\bar{{\boldsymbol{z}}}=\frac{1}{N}\sum_{i=1}^{N}{\boldsymbol{z}}_{i}$. Figure 2 shows the singular values of $C$ for SimCLR and VSimCLR. VSimCLR produces a substantially flatter spectrum, indicating a higher effective rank and thus mitigating dimensional collapse. Remarkably, on CIFAR-100, VSimCLR nearly doubles the number of dominant components compared to SimCLR. These results demonstrate that VSimCLR not only preserves semantic clustering but also leverages the embedding space more fully, and can be combined with other collapse-mitigation strategies for further gains.
4.3 Mutual Information Comparison
Figure 3: Estimate of $I({\boldsymbol{z}};{\boldsymbol{c}})$.
Figure 3 reports the estimated mutual information $I({\boldsymbol{z}};{\boldsymbol{c}})$ between the learned embeddings ${\boldsymbol{z}}$ and their true class labels ${\boldsymbol{c}}$ of CIFAR-10. We compute this using the Mixed KSG estimator gao2017estimating , which is well-suited for mixed or multimodal distributions.
Both VSimCLR and VSupCon achieve mutual information on par with—or slightly exceeding—their non-variational counterparts. In particular, during the first 200 epochs, VSimCLR exhibits lower mutual information than SimCLR, reflecting the added optimization challenge of the KL regularizer. After this initial phase, VSimCLR surpasses SimCLR and maintains higher mutual information for the remainder of training. These results indicate that VSimCLR ultimately preserves—or even improves—information between embeddings and labels, while also producing rich distributional representations.
Table 1: Classification accuracy on various datasets. We report top-1 and top-5 accuracies of SimCLR, VSimCLR, SupCon, and VSupCon across the datasets.
Method CIFAR-10 CIFAR-100 Tiny-imagenet STL10 Caltech256
Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5
SimCLR 78.42 98.52 49.56 78.84 38.95 66.89 60.44 95.80 43.14 66.15
VSimCLR 81.48 98.95 54.58 82.87 37.70 66.06 60.11 92.00 48.50 69.99
[…]
VSupCon 93.85 99.68 71.66 89.42 48.30 72.84 75.76 96.99 83.06 91.29
Figure 4: Sample images from the CIFAR-10, organized by class (columns) and sorted by their corresponding $\log{\hbox{det}}(K)$ (rows). In each column, the top image has the highest $\log{\hbox{det}}(K)$, the bottom image the lowest; the overlaid numbers indicate each image’s $\log{\hbox{det}}(K)$.
4.4 Classification
For classification, we use the posterior mean $\boldsymbol{\mu}_{\mathbf{x}}$ as the embedding and train a linear classifier. [^2] In Appendix C.2, we present additional results exploring various VSimCLR design choices.
Table 1 reports Top-1 and Top-5 accuracies on CIFAR-10, CIFAR-100, Tiny-ImageNet, STL-10, and Caltech-256. VSimCLR outperforms SimCLR on CIFAR-10 ($78.42\to 81.48$) and CIFAR-100 ($49.56\to 54.58$) in Top-1 accuracy, with similar gains in Top-5. On Caltech-256, VSimCLR also improves Top-1 accuracy substantially. Performance on Tiny-ImageNet and STL-10 remains comparable, with slight decreases (within experimental variance) likely due to the KL regularizer.
SupCon provides strong supervised baselines, and VSupCon further improves Top-1 accuracy on CIFAR-10 ($93.60\to 93.85$) and CIFAR-100 ($70.79\to 71.66$). Modest declines on Tiny-ImageNet, STL-10, and Caltech-256 reflect the trade-off of adding the KL term on datasets with higher complexity or fewer samples.
Although VCL is not explicitly designed to boost classification accuracy, both VSimCLR and VSupCon consistently match or exceed their deterministic counterparts. This demonstrates that our KL-regularized, distributional embeddings preserve the alignment and uniformity properties, while providing meaningful uncertainty estimates for downstream tasks.
4.5 Implications of Distributional Embeddings
Figure 5: Relationship between posterior dispersion and label ambiguity. Each point plots $\log{\hbox{det}}(K)$ against the entropy of human-annotated class probabilities from CIFAR-10H peterson2019human , with a first-order linear fit (red line).
We illustrate the interpretability of distributional embeddings using examples from CIFAR-10. Figure 4 displays sample images alongside the log-determinant $\log{\hbox{det}}(K)$ of their posterior covariance $K$ learned by VSimCLR. Top-row images are common class members and exhibit larger $\log{\hbox{det}}(K)$—indicating broader posterior dispersion—whereas bottom-row images are atypical or uncommon with smaller $\log{\hbox{det}}(K)$, reflecting more concentrated posteriors. [^3] Although one might assume that a larger log-determinant of a covariance matrix implies greater uncertainty (e.g., due to noise), here $\log{\hbox{det}}(K)$ measures the spread of the posterior embedding distribution. Larger values correspond to more “typical” samples with diverse latent representations, while smaller values indicate more “unique” or outlier features with tightly concentrated embeddings.
Figure 5 quantifies this effect by plotting $\log{\hbox{det}}(K)$ versus the entropy of the soft labels in the CIFAR-10H dataset peterson2019human . The negative slope of the linear regression (red line) confirms that images with lower $\log{\hbox{det}}(K)$—i.e., more concentrated posteriors—tend to have higher label entropy, and thus greater ambiguity. These results demonstrate that the dispersion of the learned posterior correlates with semantic uncertainty, highlighting the practical interpretability of VCL’s distributional embeddings. We provide additional experiments regrading the correlation between the posterior distribution and label uncertainty, and between posterior distribution and out-of-distribution samples in Appendix C.
5 Conclusion
We have introduced Variational Contrastive Learning (VCL), a decoder-free ELBO maximization framework that endows contrastive learning with principled probabilistic embeddings. By interpreting InfoNCE as a surrogate reconstruction term and regularizing via a KL divergence to a uniform prior on the unit sphere, VCL enables distributional encodings without explicit decoders. We instantiated VCL in two variants—VSimCLR and VSupCon—by replacing deterministic embeddings with samples from $q_{\theta}({\boldsymbol{z}}|{\boldsymbol{x}})$ and adding a normalized KL term.
Theoretical study and empirical experiments on CIFAR-10, CIFAR-100, Tiny-ImageNet, STL-10, and Caltech-256 demonstrate that VCL preserves the characteristic of contrastive embeddings, mitigates dimensional collapse, maintains or improves mutual information with labels, and matches or exceeds deterministic baselines in classification accuracy, all while providing meaningful posterior estimates.
Although out of scope for this paper, an intriguing future direction is to integrate VCL with generative models. Since VAEs learn embeddings by reconstructing input samples relying on decoders, and VCL learns embeddings by contrasting features, VCL with generative models may yield more generalized generative performance.
Appendix A Related work
A.1 Contrastive learning
Self-supervised contrastive learning methods [8, 46] train an encoder $f\colon\mathcal{X}\to\mathcal{S}^{d_{z}-1}$ by drawing semantically related views (positives) together in the embedding space while pushing unrelated views (negatives) apart. In the standard setup, each example is treated as its own category, and only its augmented copies count as positives. A variety of contrastive objectives—such as InfoNCE [35], Debiased Contrastive Loss [10], Unbiased Contrastive Loss [2], triplet-based losses [9, 19], and others—have been used to learn robust representations for tasks ranging from dense prediction in computer vision [54] to multimodal alignment [41, 14]. InfoNCE in particular has been shown to lower-bound mutual information [40], and subsequent work has revealed that its empirical success hinges on a balance of alignment and uniformity in the learned embeddings [47, 53]. In the supervised setting, SupCon [23] extends this idea by using class labels to define positive pairs among same-class samples, often surpassing cross-entropy training in downstream performance.
A.2 Probabilistic contrastive learning
A growing body of work has begun to integrate probabilistic latent-variable modeling with contrastive objectives. In the video domain, Park et al. represent each video clip as a Gaussian and combine them into a mixture model, learning these distributions via a stochastic contrastive loss that captures clip-level uncertainty and obviates complex augmentation schemes [36]. For 3D point clouds, Wang et al. propose a Generative Variational-Contrastive framework that models latent features as Gaussians, enforces distributional consistency across positive pairs by combining the variational autoencoder and contrastive learning [52]. In graph representation learning, Xie and Giraldo introduce Subgraph Gaussian Embedding Contrast, which maps subgraphs into a structured Gaussian space and employs optimal-transport distances for robust contrastive objectives, yielding improved classification and link-prediction performance [56].
On the theoretical front, Zimmermann et al. prove that contrastive objectives invert the data-generating process under mild conditions, uncovering a deep connection to nonlinear independent component analysis [58]. With a more generalized setting, Kirchhof et al. extend the InfoNCE loss so that the encoder predicts a full posterior distribution rather than a point, and prove that these distributions asymptotically recover the true aleatoric uncertainty of the data-generating process [25].
A.3 Variational Inference and Contrastive Learning
The most closely related line of work frames contrastive learning within a latent-variable inference paradigm via Recognition-Parametrised Models (RPMs) [1, 50]. Aitchison and Ganev introduce RPMs as a class of Bayesian models whose (unnormalized) likelihood is defined implicitly through a recognition network [1]. They show that, under RPMs, the ELBO decomposes into mutual information minus a KL term (up to a constant), and that for a suitable choice of prior the infinite-sample InfoNCE objective coincides with this ELBO. Walker et al. consider RPMs by assuming conditional independence of observations given latent variables, and develop an EM algorithm that achieves exact maximum-likelihood learning for discrete latents along with principled posterior inference [50].
Other works recast variational inference itself as a contrastive estimation task. Rhodes and Gutmann’s Variational Noise-Contrastive Estimation (VNCE) derives a variational lower bound to the standard NCE objective, enabling joint learning of model parameters and latent posteriors in unnormalized models [42]. More recently, Ward et al. propose SoftCVI, which treats VI as a classification problem: they generate “soft” pseudo-labels from the unnormalized posterior and optimize a contrastive-style objective that yields zero-variance gradients at the optimum [55].
A.4 Dimensional collapse
In contrastive self-supervised learning, several approaches have been proposed to prevent dimensional collapse by regularizing either the embedding projector or the second-order statistics of the representations. Jing et al. [22] first demonstrated that, despite the repulsive effect of negative samples, embeddings can still collapse to a low-dimensional subspace due to a combination of strong augmentations and implicit low-rank bias in weight updates. They introduced DirectCLR, which fixes a low-rank diagonal projector during training; this projector enforces the embeddings to occupy a predetermined subspace and was shown empirically to outperform SimCLR’s learned linear projector.
Following this, several works have designed novel loss functions that explicitly regularize the covariance or cross-correlation of the embedding vectors. Ermolov et al. [12] apply a whitening MSE loss so that positive pairs match under mean-square error while enforcing identity covariance. Barlow Twins [57] minimize the deviation of the normalized cross-correlation matrix from the identity, effectively performing “soft whitening” to reduce redundancy. VICReg [3] further augments this idea by combining variance, invariance, and covariance regularizers to avoid collapse without using negative samples; notably, VICReg allows its two branches to use different architectures or even modalities, enabling joint embedding across data types. More recently, He et al. [16] showed that orthogonal regularization of encoder weight matrices preserves representation diversity and prevents collapse.
Appendix B Proofs
B.1 Proof of Proposition 3.1
Proof.
Optimal critic [31] for InfoNCE satisfies that
$\psi^{\star}({\boldsymbol{x}},{\boldsymbol{z}})$ $\propto\log\frac{p({\boldsymbol{x}}|{\boldsymbol{z}})}{p({\boldsymbol{x}})}+\alpha({\boldsymbol{z}}),$ (25)
where $\alpha({\boldsymbol{z}})$ only depends on ${\boldsymbol{z}}$. With the optimal critic, we then have
$I_{\rm NCE}({\boldsymbol{x}};{\boldsymbol{x}}^{\prime})$ $=-\mathbb{E}\left[\log\frac{e^{\psi({\boldsymbol{z}},{\boldsymbol{z}}^{\prime}_{i})}}{\sum_{j=1}^{N}e^{\psi({\boldsymbol{z}},{\boldsymbol{z}}^{\prime}_{j})}}\right]$
$=-\mathbb{E}\left[\log\frac{p({\boldsymbol{z}}|{\boldsymbol{z}}^{\prime}_{i})}{\sum_{j=1}^{N}p({\boldsymbol{z}}|{\boldsymbol{z}}^{\prime}_{j})}\right]$
$=-\mathbb{E}\left[\log\frac{p({\boldsymbol{z}}|{\boldsymbol{z}}^{\prime}_{i})}{\frac{1}{N}\sum_{j=1}^{N}p({\boldsymbol{z}}|{\boldsymbol{z}}^{\prime}_{j})}\right]+\log N.$ (26)
Rearranging (25), we obtain
$-I_{\rm NCE}({\boldsymbol{x}};{\boldsymbol{x}}^{\prime})+\log N$
$=\mathbb{E}\left[\log\frac{p({\boldsymbol{z}}|{\boldsymbol{z}}^{\prime}_{i})}{\frac{1}{N}\sum_{j=1}^{N}p({\boldsymbol{z}}|{\boldsymbol{z}}^{\prime}_{j})}\right]$
=$\overset{N\to\infty}{=}\mathbb{E}\left[\log\frac{p({\boldsymbol{z}}|{\boldsymbol{z}}^{\prime}_{i})}{\mathbb{E}_{p({\boldsymbol{z}}^{\prime})}[p({\boldsymbol{z}}|{\boldsymbol{z}}^{\prime})]}\right]$
$=\mathbb{E}\left[\log\frac{p({\boldsymbol{z}}|{\boldsymbol{z}}^{\prime}_{i})}{p({\boldsymbol{z}})}\right]$
$=\mathbb{E}\left[\log\frac{p({\boldsymbol{z}}^{\prime}_{i}|{\boldsymbol{z}})}{p({\boldsymbol{z}}^{\prime}_{i})}\right]$
$=\mathbb{E}_{q_{\theta}({\boldsymbol{z}}_{i}^{\prime}|{\boldsymbol{x}})q_{\theta}({\boldsymbol{z}}|{\boldsymbol{x}})}\left[\log p({\boldsymbol{z}}^{\prime}_{i}|{\boldsymbol{z}})\right]+\mathbb{E}_{q_{\theta}({\boldsymbol{z}}_{i}^{\prime}|{\boldsymbol{x}})}\left[\log p({\boldsymbol{z}}^{\prime}_{i})\right]$
$=\mathbb{E}_{q_{\theta}({\boldsymbol{z}}_{i}^{\prime}|{\boldsymbol{x}})q_{\theta}({\boldsymbol{z}}|{\boldsymbol{x}})}\left[\log p({\boldsymbol{z}}^{\prime}_{i}|{\boldsymbol{z}})\right]+\mathbb{E}_{q_{\theta}({\boldsymbol{z}}_{i}^{\prime}|{\boldsymbol{x}})}\left[\log\frac{p({\boldsymbol{z}}^{\prime}_{i})}{q_{\theta}({\boldsymbol{z}}_{i}^{\prime}|{\boldsymbol{x}})}\right]+\mathbb{E}_{q_{\theta}({\boldsymbol{z}}_{i}^{\prime}|{\boldsymbol{x}})}\left[\log q_{\theta}({\boldsymbol{z}}^{\prime}_{i}|{\boldsymbol{x}})\right]$
$=\mathbb{E}_{q_{\theta}({\boldsymbol{z}}_{i}^{\prime}|{\boldsymbol{x}})q_{\theta}({\boldsymbol{z}}|{\boldsymbol{x}})}\left[\log p({\boldsymbol{z}}^{\prime}_{i}|{\boldsymbol{z}})\right]-D(q_{\theta}({\boldsymbol{z}}_{i}^{\prime}|{\boldsymbol{x}})\|p({\boldsymbol{z}}^{\prime}_{i}))-H(q_{\theta}({\boldsymbol{z}}^{\prime}_{i}|{\boldsymbol{x}})),$ (27)
where the second equality follows by the strong law of large numbers. Rewriting ${\boldsymbol{z}}_{i}^{\prime}$ with ${\boldsymbol{z}}^{\prime}$, this concludes the proof of Proposition 3.1
B.2 Formal statement of Theorem 3.2 and its proof
In this section, we provide detailed problem setup, formal statement of Theorem 3.2, its extension to random augmentation, and the proof of Theorem 3.2.
B.2.1 Problem setup and formal statement of Theorem 3.2
First, we introduce the class of the deep forward neural network with bounded weights and Lipschitz activations.
Let $\{\bm{x}_{i}\}_{i=1}^{N}\subset\mathcal{X}$ be independent and identically distributed samples from a distribution $\mathcal{D}$ on $\mathcal{X}$. We consider depth-$L$ feed-forward networks $f_{\bm{\Theta}}:\mathcal{X}\to\mathbb{R}^{2d}$ of the form
$f_{\bm{\Theta}}(\bm{x})=\alpha_{L}\Bigl{(}\bm{\Theta}^{L}\,\alpha_{L-1}\bigl{(}\bm{\Theta}^{L-1}\,\cdots\alpha_{1}(\bm{\Theta}^{1}\bm{x})\bigr{)}\Bigr{)},$
where each activation $\alpha_{\ell}:\mathbb{R}^{d_{\ell}}\to\mathbb{R}^{d_{\ell}}$ is $\sigma_{\ell}$-Lipschitz. For each layer $\ell=1,\dots,L$, the weight matrices lie in
$\bm{\Xi}_{\ell}:=\bigl{\{}\bm{\Theta}^{\ell}\in\mathbb{R}^{d_{\ell}\times d_{\ell-1}}:\|\bm{\Theta}^{\ell}\|_{2}\leq\rho_{\ell},\;\|(\bm{\Theta}^{\ell}-\bm{\Theta}_{0}^{\ell})^{\top}\|_{2,1}\leq a_{\ell}\bigr{\}},$ (28)
where $\|\cdot\|_{2}$ is the spectral norm and $\|\cdot\|_{2,1}$ is the sum of column-wise $\ell_{2}$ norms, and $\{\bm{\Theta}_{0}^{\ell}\}_{\ell=1}^{L}$ are fixed reference matrices. Set $\bm{\Xi}=\prod_{\ell=1}^{L}\bm{\Xi}_{\ell}$ and define
$\mathcal{F}_{\bm{\Xi}}=\bigl{\{}f_{\bm{\Theta}}:\bm{\Theta}\in\bm{\Xi}\bigr{\}}.$ (29)
In VCL the encoder output splits as
$f_{\bm{\Theta}}(\bm{x})=\begin{pmatrix}\mu_{\bm{\Theta}}(\bm{x})\\ \varepsilon_{\bm{\Theta}}(\bm{x})\end{pmatrix},\quad\mu_{\bm{\Theta}},\,\varepsilon_{\bm{\Theta}}:\mathcal{X}\to\mathbb{R}^{d}.$ (30)
Given $\bm{x}$, the KL regularizer by output of $f_{\bm{\Theta}}(\bm{x})$ in (30) is
$D\bigl{(}f_{\bm{\Theta}};\bm{x}\bigr{)}=\frac{1}{2d}\Bigl{(}\|\mu_{\bm{\Theta}}(\bm{x})\|_{2}^{2}+\bigl{\|}\exp(\varepsilon_{\bm{\Theta}}(\bm{x}))\bigr{\|}_{1}-d-\bigl{\langle}\bm{1}_{d},\varepsilon_{\bm{\Theta}}(\bm{x})\bigr{\rangle}\Bigr{)}.$ (31)
Theorem B.1.
Fix $\delta\in(0,1)$. Under the assumptions on $\{\bm{x}_{i}\}_{i=1}^{N}$ and definitions on $\mathcal{F}_{\bm{\Xi}}$ (29) and $D\bigl{(}f_{\bm{\Theta}};\bm{x}\bigr{)}$ (31), it holds that
$\sup_{f_{\bm{\Theta}}\in\mathcal{F}_{\bm{\Xi}}}\Bigl{[}\mathbb{E}_{\bm{x}\sim\mathcal{D}}\,D(f_{\bm{\Theta}};\bm{x})$ $-\frac{1}{N}\sum_{i=1}^{N}D(f_{\bm{\Theta}};\bm{x}_{i})\Bigr{]}$
$\leq\widetilde{\mathcal{O}}\!\Bigl{(}\frac{\eta\,B_{x}\prod_{\ell=1}^{L}(\rho_{\ell}\,\sigma_{\ell})}{\sqrt{N}}\Bigl{[}\sqrt{\log W}\,\Bigl{(}\sum_{\ell=1}^{L}\bigl{(}a_{\ell}/\rho_{\ell}\bigr{)}^{2/3}\Bigr{)}^{3/2}\;\vee\;\sqrt{\log(1/\delta)}\Bigr{]}\Bigr{)},$
with probability at least $1-\delta$, where
$W=\max_{\ell\in[L]}d_{\ell},\quad B_{x}=\sup_{\bm{x}\in\mathcal{X}}\|\bm{x}\|_{2},\quad\eta=\mathcal{O}\!\Bigl{(}1\vee e^{B_{x}\prod_{\ell=1}^{L}\rho_{\ell}\,\sigma_{\ell}}\Bigr{)}.$
Random augmentations.
One might be concerned that the encoder in (30) uses raw inputs, whereas in practice each sample $\bm{x}_{i}$ is fed through two random augmentations $t^{\prime}_{i},t^{\prime\prime}_{i}\sim\mathcal{T}$ (see Section 2.1). To deal with this concern, we show that exactly the same bound of Theorem B.1 holds under this more practical setting.
Let $\mathcal{T}$ be a distribution over augmentation maps $t:\mathcal{X}\to\mathcal{X}$. Define the augmented data law
$\mathcal{D}_{\rm aug}\;:=\;\operatorname{Law}_{\,\bm{x}\sim\mathcal{D},\;t\sim\mathcal{T}}\bigl{(}t(\bm{x})\bigr{)}\;=\;\mathcal{T}_{\#}\mathcal{D},$
where $\mathcal{T}_{\#}\mathcal{D}$ denotes the push-forward of the measure $\mathcal{D}$ through the (random) map distribution $\mathcal{T}$.
If
$t_{i}\;\overset{\mathrm{i.i.d.}}{\sim}\;\mathcal{T},\quad\bm{x}_{i}^{\prime}=t_{i}(\bm{x}_{i}),$
then by construction
$\bm{x}_{1}^{\prime},\dots,\bm{x}_{N}^{\prime}\;\overset{\mathrm{i.i.d.}}{\sim}\;\mathcal{D}_{\rm aug},$
so the conditions of Theorem B.1 remain satisfied with $\mathcal{D}_{\rm aug}$ in place of $\mathcal{D}$.
Corollary B.2.
Let $\{t^{\prime}_{i}\}_{i=1}^{N}$ and $\{t^{\prime\prime}_{i}\}_{i=1}^{N}$ be independent samples from $\mathcal{T}$, and set $\bm{x}_{i}^{\prime}=t^{\prime}_{i}(\bm{x}_{i})$, $\bm{x}_{i}^{\prime\prime}=t^{\prime\prime}_{i}(\bm{x}_{i})$. Under the assumptions of Theorem B.1, with probability at least $1-\delta$ we have
$\sup_{f_{\bm{\Theta}}\in\mathcal{F}_{\bm{\Xi}}}\Bigl{[}\mathbb{E}_{\bm{x}\sim\mathcal{D},\,t\sim\mathcal{T}}$ $D\bigl{(}f_{\bm{\Theta}};t(\bm{x})\bigr{)}-\frac{1}{2N}\sum_{i=1}^{N}\bigl{(}D(f_{\bm{\Theta}};\bm{x}_{i}^{\prime})+D(f_{\bm{\Theta}};\bm{x}_{i}^{\prime\prime})\bigr{)}\Bigr{]}$
$\leq\widetilde{\mathcal{O}}\!\Bigl{(}\frac{\eta\,B_{x,t}\prod_{\ell=1}^{L}(\rho_{\ell}\sigma_{\ell})}{\sqrt{N}}\Bigl{[}\sqrt{\log W}\,\Bigl{(}\sum_{\ell=1}^{L}(a_{\ell}/\rho_{\ell})^{2/3}\Bigr{)}^{3/2}\;\vee\;\sqrt{\log(1/\delta)}\Bigr{]}\Bigr{)},$
where $B_{x,t}=\sup_{\bm{x}\in\mathcal{X},\,t\in\mathcal{T}}\|t(\bm{x})\|_{2}$.
One might wonder whether $B_{x,t}$ is finite in Corollary B.2. In practice, common vision augmentations—such as random resized cropping and horizontal flips [27, 37], color jittering (brightness, contrast, saturation perturbations) [44, 8], and additive Gaussian blur or noise [18]—always produce outputs that remain within a bounded neighborhood of the original inputs. For example, cropping a $256\times 256$ image to $224\times 224$ keeps all pixel values in their original range, and color jitter is applied with limited intensity so that augmented images lie close in Euclidean norm to the source. Even when combining multiple operations (as in AugMix), the convex mixtures ensure augmented samples do not “drift” outside the natural image manifold. Hence, $\sup_{\bm{x}\in\operatorname{\mathcal{X}},\,t\in\mathcal{T}}\|t(\bm{x})\|_{2}$ remains finite under these widely adopted schemes.
Before proving Theorem B.1, we need to introduce some background on Rademacher complexity, Dudley’s entropy-integral bound, and Lipschitz of KL divergence, which are necessary in the proof.
B.2.2 Backround: Rademacher complexity and Dudley’s entropy-integral bound
We introduce the background on Rademacher complexity, Dudley’s entropy-integral bound, and covering number for the function class $\mathcal{F}_{\bm{\Xi}}$ in (29), which are used in the proof of Theorem B.1.
Lemma B.3 (Rademacher generalization bound).
[33] Let $\mathcal{X}$ be a vector space and $\mathcal{D}$ a distribution over $\mathcal{X}$. Let $\mathcal{F}$ be a class of functions $f:\mathcal{X}\to[a,b]$. Draw samples $\mathcal{S}=\{\bm{x}_{i}\}_{i=1}^{N}\overset{\mathrm{i.i.d}}{\sim}\mathcal{D}^{n}$, and define the empirical Rademacher complexity
^$\hat{\mathfrak{R}}_{S}(\mathcal{F})\;=\;\mathbb{E}_{\sigma}\Biggl{[}\sup_{f\in\mathcal{F}}\frac{1}{N}\sum_{i=1}^{N}t_{i}\,f(\bm{x}_{i})\Biggr{]},$ (32)
where $t_{i}\in\{-1,1\}$ are independent Rademacher variables. Then for any $\delta\in(0,1)$, with probability at least $1-\delta$,
[…]
Lemma B.4 (Dudley’s entropy-integral bound).
[4, Lemma 8.5] Let $\mathcal{X}$ be a vector space and $\mathcal{F}$ a class of functions $f:\mathcal{X}\to\mathbb{R}$. For a sample $\mathcal{S}=\{\bm{x}_{i}\}_{i=1}^{N}$, define
$\|f\|_{L_{2}(\mathcal{S})}\;=\;\left(\tfrac{1}{N}\sum_{i=1}^{N}f(\bm{x}_{i})^{2}\right)^{1/2},\quad B_{\mathcal{F}}\;=\;\sup_{f\in\mathcal{F}}\|f\|_{L_{2}(\mathcal{S})}.$
Then the empirical Rademacher complexity of $\mathcal{F}$ (32) satisfies
[…]
where $\mathcal{N}(\mathcal{F},\varepsilon,L_{2}(\mathcal{S}))$ is the covering number of $\mathcal{F}$ under the $\|\cdot\|_{L_{2}(\mathcal{S})}$ metric.
Lemma B.5 (Covering-number bound for $\mathcal{F}_{\bm{\Xi}}$).
[21, Proposition 4] Let $\mathcal{F}_{\bm{\Xi}}$ be the class of $L$-layer neural networks with bounded weights and Lipschitz activation functions, in (29). Let
$\mathcal{S}=\{\bm{x}_{i}\}_{i=1}^{N}\subset\mathcal{X}$
be any fixed dataset. Define
¯$\overline{\mathcal{R}}_{\Xi}^{2/3}=\sum_{l=1}^{L}\bigl{(}a_{l}\,B_{l-1}\,\rho_{l+}\bigr{)}^{2/3},\quad$
$\rho_{l+}=\rho_{l}\;\prod_{m=l+1}^{L}\bigl{(}\rho_{m}\,s_{m}\bigr{)},$
$B_{l}=\sup_{\bm{x}\in\mathcal{X}}\,\sup_{\bm{\Theta}\in\bm{\Xi}}\bigl{\|}f_{\bm{\Theta}}^{1\to l}(\bm{x})\bigr{\|}_{2},$
where $f_{\bm{\Theta}}^{1\to l}(\bm{x})$ is the output of the first $l$ layers of the network. Then for every $\varepsilon>0$,
$\log\mathcal{N}\!\bigl{(}\mathcal{F}_{\bm{\Xi}},\varepsilon,L_{\infty,2}(S)\bigr{)}\;\leq\;\frac{64\,\overline{\mathcal{R}}_{\Xi}^{2}}{\varepsilon^{2}}\;\log\!\Bigl{(}\bigl{(}\tfrac{11\,\overline{\mathcal{R}}_{\Xi}}{\varepsilon}+7\bigr{)}\,N\,W\Bigr{)}.$
B.2.3 $\ell_{2}$-Lipschitzness of KL divergence
We introduce the Lipschitzness of the KL divergence term $D(f_{\bm{\Theta}};\bm{x})$ in (31) with respect to the $\ell_{2}$ norm when we assume the output of the encoder is bounded by $B_{L}$. The Lipschitzness of the KL divergence term is crucial for the proof of Theorem B.1. We will specify the bound $B_{L}$ later in the proof of Theorem B.1.
Lemma B.6 ($\ell_{2}$–Lipschitzness of the KL term).
For some constant $B_{L}>0$, define
$V_{B_{L}}\coloneqq\left\{[\bm{\mu};\bm{\varepsilon}]\in\mathbb{R}^{2d}:\bm{\mu}\in\mathbb{R}^{d},\bm{\varepsilon}\in\mathbb{R}^{d},\|[\bm{\mu};\bm{\varepsilon}]\|_{2}\leq B_{L}.\right\},$
and let
$D([\bm{\mu};\bm{\varepsilon}]):=\frac{1}{2d}\sum_{j=1}^{d}\Bigl{(}\bm{\mu}_{j}^{2}+\exp(\bm{\varepsilon}_{j})-1-\bm{\varepsilon}_{j}\Bigr{)}.$ (33)
Then, for any two pairs $[\bm{\mu};\bm{\varepsilon}],[\bm{\mu}^{\prime};\bm{\varepsilon}^{\prime}]\in V_{B_{L}}$, it holds that
$\bigl{|}D([\bm{\mu};\bm{\varepsilon}])-D([\bm{\mu}^{\prime};\bm{\varepsilon}^{\prime}])\bigr{|}\;\leq\;\;\frac{\eta}{\sqrt{d}}\cdot\bigl{\|}([\bm{\mu};\bm{\varepsilon}])-([\bm{\mu}^{\prime};\bm{\varepsilon}^{\prime}])\bigr{\|}_{2},$
where $\eta=B_{L}+(1+\exp(B_{L}))/2.$
Proof.
Compute the gradient of $D([\bm{\mu};\bm{\varepsilon}])$ in (33) as
$\nabla D\bigl{(}[\bm{\mu};\bm{\varepsilon}]\bigr{)}=\frac{1}{d}\begin{bmatrix}\bm{\mu}\\ \frac{\exp(\bm{\varepsilon})-\bm{1}_{d}}{2}\end{bmatrix}$ (34)
where $\bm{1}_{d}\coloneqq[1,\dots,1]\in\mathbb{R}^{d}$.
Fix $[\bm{\mu};\bm{\varepsilon}],[\bm{\mu}^{\prime};\bm{\varepsilon}^{\prime}]\in V_{B_{L}}$ arbitrarily. Then, applying Cauchy–Schwarz inequality with the convexity of $D([\bm{\mu};\bm{\varepsilon}])$ yields
$\bigl{|}D([\bm{\mu};\bm{\varepsilon}])-D([\bm{\mu}^{\prime};\bm{\varepsilon}^{\prime}])\bigr{|}\;\leq\;\nabla D([\bm{\mu};\bm{\varepsilon}])^{\top}([\bm{\mu};\bm{\varepsilon}]-[\bm{\mu}^{\prime};\bm{\varepsilon}^{\prime}])\;\leq\;\bigl{\|}\nabla D([\bm{\mu};\bm{\varepsilon}])\bigr{\|}_{2}\;\|[\bm{\mu};\bm{\varepsilon}]-[\bm{\mu}^{\prime};\bm{\varepsilon}^{\prime}]\|_{2},$
Hence, it suffices to show $\bigl{\|}\nabla D([\bm{\mu};\bm{\varepsilon}])\bigr{\|}_{2}\leq\eta/\sqrt{d}$ for all $[\bm{\mu};\bm{\varepsilon}]\in V_{B_{L}}$. By (34), triangle inequality gives
$\sup_{[\bm{\mu};\bm{\varepsilon}]\in V_{B_{L}}}\|\nabla D([\bm{\mu};\bm{\varepsilon}])\|_{2}$ $\leq\frac{1}{d}\sup_{[\bm{\mu};\bm{\varepsilon}]\in V_{B_{L}}}\|\bm{\mu}\|_{2}+\frac{1}{2d}\sup_{[\bm{\mu};\bm{\varepsilon}]\in V_{B_{L}}}\|\exp(\bm{\varepsilon})\|_{2}+\frac{1}{2\sqrt{d}}$
${\leq}\frac{1}{d}B_{L}+\frac{1}{2\sqrt{d}}\exp({B_{L}})+\frac{1}{2\sqrt{d}}$
where in the last inequality, we use $\|\bm{\mu}\|_{2}\leq\|[\bm{\mu};\bm{\varepsilon}]\|_{2}\leq B_{L}$ and
$\|\exp(\bm{\varepsilon})\|_{2}=\sqrt{\sum_{j=1}^{d}\exp(2\bm{\varepsilon}_{j})}\leq\sqrt{d}\exp(B_{L}).$
As $\frac{1}{\sqrt{d}}\leq 1$ for any $d\geq 1$, we have shown $\bigl{\|}\nabla D([\bm{\mu};\bm{\varepsilon}])\bigr{\|}_{2}\leq\eta/\sqrt{d}$. ∎
B.2.4 Proof of Theorem B.1
Note that $\rho_{\ell}$ bounded spectral norm of weight matrices and $\sigma_{\ell}$-Lipschitzness of activation function for each layer $\ell\in[L]$ yield
$\sup_{f_{\bm{\Theta}}\in\mathcal{F}_{\bm{\Xi}},\bm{x}\in\mathcal{X}}\|f_{\bm{\Theta}}(\bm{x})\|_{2}\leq\underbrace{\sup_{\bm{x}\in\mathcal{X}}\left\|\bm{x}\right\|_{2}\cdot\prod_{\ell=1}^{L}\rho_{\ell}\sigma_{\ell}}_{B_{L}}.$ (35)
Therefore, for all $\bm{x}\in\mathcal{X}$, $\left\|[{\bm{\mu}}_{\bm{\Theta}}(\bm{x});{\bm{\varepsilon}}_{\bm{\Theta}}(\bm{x})]\right\|_{2}\leq B_{L}$ holds for all $\bm{\Theta}\in\mathcal{F}_{\bm{\Xi}}$. Then, we can apply the $\ell_{2}$ Lipschitzness of KL divergence in Lemma B.6 to bound the KL divergence term $D(f_{\bm{\Theta}};\bm{x})$ in (31) as
$\left|D(f_{\bm{\Theta}};\bm{x}_{i})-D(f_{\bm{\Theta}^{\prime}};\bm{x}_{i})\right|$ $\leq\eta\cdot\left\|f_{\bm{\Theta}}(\bm{x}_{i})-f_{\bm{\Theta}^{\prime}}(\bm{x}_{i})\right\|_{2}$ (36)
$\leq\eta\cdot\max_{i\in[N]}\left\|f_{\bm{\Theta}}(\bm{x}_{i})-f_{\bm{\Theta}^{\prime}}(\bm{x}_{i})\right\|_{2}$
$=\frac{\eta}{\sqrt{d}}\cdot\|f_{\bm{\Theta}}-f_{\bm{\Theta}^{\prime}}\|_{L_{\infty,2(\mathcal{S})}},\quad\quad\forall f_{\bm{\Theta}},f_{\bm{\Theta}^{\prime}}\in\mathcal{F}_{\bm{\Xi}}.$
Since the bound in (36) yields
$\left\|D(f_{\bm{\Theta}})-D(f_{\bm{\Theta}^{\prime}})\right\|_{L_{2}(\mathcal{S})}$ $=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(D(f_{\bm{\Theta};\bm{x}_{i}})-D(f_{\bm{\Theta}^{\prime};\bm{x}_{i}})\right)^{2}}$
$\leq\frac{\eta}{\sqrt{d}}\cdot\|f_{\bm{\Theta}}-f_{\bm{\Theta}^{\prime}}\|_{L_{\infty,2(\mathcal{S})}},\quad\forall f_{\bm{\Theta}},f_{\bm{\Theta}^{\prime}}\in\mathcal{F}_{\bm{\Xi}},$
we have
$\log\mathcal{N}\left(\Delta,\epsilon,{L_{2}}(\mathcal{S})\right)\leq\log\mathcal{N}\left(\mathcal{F}_{\bm{\Xi}},\frac{\epsilon\sqrt{d}}{\eta},L_{\infty,2}(\mathcal{S})\right),$
where we define $\Delta$ as the class of KL divergence terms $D(f_{\bm{\Theta}})$ in (31), i.e.,
$\Delta:=\left\{D(f_{\bm{\Theta}}):f_{\bm{\Theta}}\in\mathcal{F}_{\bm{\Xi}}\right\}.$
Now, we can apply the covering number bound in Lemma B.5 to bound the covering number $\mathcal{N}\left(\Delta,\epsilon,{L_{2}}(\mathcal{S})\right)$ as
$\mathcal{N}\!\bigl{(}\Delta,\epsilon,L_{2}(\mathcal{S})\bigr{)}\;\leq\;\frac{64\,\eta^{2}\,\overline{\mathcal{R}}_{\bm{\Xi}}^{2}}{\epsilon^{2}d}\;\log\!\Bigl{(}\bigl{(}\tfrac{11\eta\overline{\mathcal{R}}_{\bm{\Xi}}}{\epsilon\sqrt{d}}+7\bigr{)}\,N\,W\Bigr{)}.$ (37)
Define $B_{\Delta}:=\sup_{D(f_{\bm{\Theta}})\in\Delta}\|D(f_{\bm{\Theta}})\|_{L_{2}(\mathcal{S})}$. We further bound the empirical Rademacher complexity $\hat{R}_{\mathcal{S}}(\Delta)$ by applying the Dudley’s entropy integral Lemma B.4 with the bound on covering number (37) by the choosing $\alpha=\frac{1}{\sqrt{N}}$:
[…]
(37)${\leq}\frac{4}{\sqrt{N}}+\frac{96\eta\overline{\mathcal{R}}_{\bm{\Xi}}}{\sqrt{Nd}}\sqrt{\log\left(\left(\frac{11\eta\overline{\mathcal{R}}_{\bm{\Xi}}\sqrt{N}}{\sqrt{d}}+7\right)NW\right)}\log(\sqrt{N}B_{\Delta}).$ ) .
Now, we apply the Rademacher generalization bound Lemma B.3 with the bound on the empirical Rademacher complexity in (38) by letting $M=\sup_{\bm{\Theta}\in\bm{\Xi},\bm{x}\in\mathcal{X}}\left|D(f_{\bm{\Theta}};\bm{x})\right|$. Hence, it holds that
$\sup_{f_{\bm{\Theta}}\in\mathcal{F}_{\bm{\Xi}}}\left|\operatorname{\mathbb{E}}_{\bm{x}\sim\mathcal{D}}\left[D(f_{\bm{\Theta}};\bm{x})\right]-\frac{1}{N}\sum_{i=1}^{N}D(f_{\bm{\Theta}}(\bm{x}_{i}))\right|$ (39)
$\quad\leq 2{\widehat{\mathfrak{R}}_{\mathcal{S}}(\Delta)}+3M\sqrt{\frac{\log{2}/{\delta}}{2N}}$
$\quad\leq\frac{8}{\sqrt{N}}+\frac{192\eta\overline{\mathcal{R}}_{\bm{\Xi}}}{\sqrt{Nd}}\sqrt{\log\left(\left(\frac{11\eta\overline{\mathcal{R}}_{\bm{\Xi}}\sqrt{N}}{\sqrt{d}}+7\right)NW\right)}\log(\sqrt{N}B_{\Delta})+3M\sqrt{\frac{\log{2}/{\delta}}{2N}},$
with probability at least $1-\delta$.
Suppose that the remainder of the proof is conditioned on the event (39). We will now show the upper bounds on $B_{\Delta}$, $M$, and ¯$\operatorname{\overline{\mathcal{R}}_{\bm{\Xi}}}$ in (39).
Bound on $B_{{\Delta}}$:
Using $\ell_{2}$ lipschitzness of $D(f_{\bm{\Theta}};\bm{x}_{i})$ in Lemma B.6 gives
$\left|D(f_{\bm{\Theta}};\bm{x}_{i})\right|\leq\eta\|f_{\bm{\Theta}}(\bm{x}_{i})\|_{2},\quad\forall f_{\bm{\Theta}}\in\mathcal{F}_{\bm{\Xi}},\quad\forall\bm{x}_{i}\in\mathcal{S}.$
This directly gives us
$B_{\Delta}=\sup_{D(f_{\bm{\Theta}})\in\Delta}\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(D(f_{\bm{\Theta}};\bm{x}_{i})\right)^{2}}$ $\leq\sqrt{\frac{\eta^{2}}{N}\sum_{i=1}^{N}\sup_{\bm{\Theta}\in\bm{\Xi}}\|f_{\bm{\Theta}}(\bm{x}_{i})\|_{2}^{2}}$ (40)
$\leq\eta B_{L},$
where $B_{L}$ is defined in (35).
Bound on $M$:
Upper bound on $M$ can be obtained in a similar way as the bound on $B_{\Delta}$ in (40):
$M=\sup_{\bm{\Theta}\in\bm{\Xi},\bm{x}\in\mathcal{X}}|D(f_{\bm{\Theta}};\bm{x})|\leq\eta\sup_{\bm{\Theta}\in\bm{\Xi},\bm{x}\in\mathcal{X}}\|f_{\bm{\Theta}}(\bm{x})\|_{2}\leq\eta B_{L}.$ (41)
Bound on ¯$\operatorname{\overline{\mathcal{R}}_{\bm{\Xi}}}$:
In the proof of [21, Theorem 1], Hieu et al. show an upper bound on ¯$\operatorname{\overline{\mathcal{R}}_{\bm{\Xi}}}$ which is
¯$\operatorname{\overline{\mathcal{R}}_{\bm{\Xi}}}\leq\sup_{\bm{x}\in\mathcal{X}}\|\bm{x}\|_{2}\prod_{\ell=1}^{L}\sigma_{\ell}\rho_{\ell}\left(\sum_{\ell^{\prime}=1}^{L}(a_{\ell^{\prime}}/\rho_{\ell^{\prime}})^{2/3}\right)^{3/2}.$ (42)
For the sake of simplicity, we denote $B_{x}=\sup_{\bm{x}\in\mathcal{X}}\|\bm{x}\|_{2}$. Putting the bounds (40), (41), and (42) into (39) provides
$\sup_{f_{\bm{\Theta}}\in\mathcal{F}_{\bm{\Xi}}}\left|\operatorname{\mathbb{E}}_{\bm{x}\sim\mathcal{D}}\left[D(f_{\bm{\Theta}};\bm{x})\right]-\frac{1}{N}\sum_{i=1}^{N}D(f_{\bm{\Theta}}(\bm{x}_{i}))\right|$
$\leq\tilde{\mathcal{O}}\left(\frac{\eta}{\sqrt{N}}\sqrt{\log(W)}B_{x}\prod_{\ell=1}^{L}\sigma_{\ell}\rho_{\ell}\left(\sum_{\ell^{\prime}=1}^{L}(a_{\ell^{\prime}}/\rho_{\ell^{\prime}})^{3/2}\right)^{{3}/{2}}+\eta B_{x}\prod_{\ell=1}^{L}(\rho_{\ell}\sigma_{\ell})\sqrt{\frac{\log(1/\delta)}{N}}\right)$
$\leq\tilde{\mathcal{O}}\left(\frac{\eta B_{x}\prod_{\ell=1}^{L}\sigma_{\ell}\rho_{\ell}}{\sqrt{N}}\left[\sqrt{\log(W)}\left(\sum_{\ell^{\prime}=1}^{L}(a_{\ell^{\prime}}/\rho_{\ell^{\prime}})^{3/2}\right)^{{3}/{2}}\vee\sqrt{\log(1/\delta)}\right]\right).$
Appendix C Experiments
C.1 Implications of Distributional Embeddings
Distributional (probabilistic) embeddings offer several useful properties, such as quantifying uncertainty and defining probabilistic distances between samples or classes.
Figure 6: Relationship between posterior dispersion and label ambiguity. Each point plots the trace of $K$ (${\hbox{tr}}(K)$) against the entropy of human-annotated class probabilities from CIFAR-10H [38], with a first-order linear fit (red line). Similar to the result in Figurer 5, the dispersion is negatively correlated with label ambiguity.
As shown in Figure 4, different samples exhibit varying degrees of posterior dispersion, which can serve as an uncertainty measure. In addition to comparing $\log{\hbox{det}}(K)$ with label entropy, we also compare the trace of $K$ (denoted by $\mathrm{tr}(K)$) against label entropy in Figure 6. Once again, these two quantities are negatively correlated under a first-order linear fit. This indicates that VSimCLR assigns higher posterior dispersion to inputs with greater label uncertainty. Conversely, inputs that humans classify unambiguously—i.e., prototypical class examples—receive posteriors with larger dispersion, suggesting that their latent representations span a broader region of the class-specific embedding space. In contrast, ambiguous or outlier inputs yield posteriors with smaller dispersion, reflecting more concentrated latent distributions. Therefore, posterior dispersion can be effectively used as a measure of uncertainty.
In Figure LABEL:fig:mu_K, we compare the class-wise averages of the posterior parameters $\|\boldsymbol{\mu}\|$ and $\log{\hbox{det}}(K)$ for VSimCLR and VSupCon. The results indicate that different classes exhibit different levels of dispersion. Notably, VSimCLR and VSupCon produce somewhat similar patterns in both $\|\boldsymbol{\mu}\|$ and $\log{\hbox{det}}(K)$, despite being trained independently. For instance, the cat and dog classes show lower values of $\|\boldsymbol{\mu}\|$ and $\log{\hbox{det}}(K)$ under both methods, whereas the truck class attains the highest $\|\boldsymbol{\mu}\|$. Table 2 provides the detailed $\log{\hbox{det}}(K)$ values for each class.
We also compare the average posterior parameters learned by VSimCLR on in-distribution (CIFAR-10) versus out-of-distribution (SVHN [34]) samples. In the experiment shown in Figure LABEL:fig:ood, VSimCLR is trained on the CIFAR-10 training set, and we then extract the posterior parameters $\boldsymbol{\mu}$ and $K$ on the CIFAR-10 and SVHN test sets. Figure LABEL:fig:ood plots these parameters for each dataset, with black markers denoting the class-averaged values. While the mean parameter points are similar between CIFAR-10 and SVHN, the SVHN samples exhibit substantially greater dispersion in their posterior parameter distribution.
Table 2: Log-determinant of average posterior covariance $K$ for each CIFAR-10 class.
Index Class $\log{\hbox{det}}(K)$
0 airplane -182.207
1 automobile -181.691
2 bird -183.713
3 cat -191.317
4 deer -184.969
5 dog -185.432
6 frog -182.125
7 horse -179.331
8 ship -185.991
9 truck -188.179
C.2 Distributional Contrastive Loss
In addition to the contrastive loss on embeddings, it is worthwhile to contrast the posterior distributions within the VCL framework. Specifically, we aim to pull together the posteriors corresponding to different augmentations of the same input and to push apart posteriors from distinct inputs. To incorporate this into VCL, we introduce the DistNCE loss, a contrastive objective over posterior parameters, defined as
$D_{\mathrm{DistNCE}}(\theta)$ $=-\mathbb{E}\left[\log\frac{\exp\bigl{(}s(\theta,\theta^{+})\bigr{)}}{\displaystyle\sum_{j}\exp\bigl{(}s(\theta,\theta_{j})\bigr{)}}\right],$ (43)
where $\theta$ denotes the posterior parameters $(\boldsymbol{\mu},K)$, $\theta^{+}$ is the positive-pair parameter for the same input, and $\{\theta_{j}\}_{j\neq+}$ are negative-pair parameters from other inputs. The expectation is taken over the joint distribution $p(\theta,\theta^{+})\prod_{j\neq+}p(\theta_{j})$.
Moreover, we increase the number of posterior samples used for the InfoNCE loss. Specifically, we draw $m$ samples $\{\mathbf{z}^{(k)}\}_{k=1}^{m}$ from each posterior, resulting in an $m$-fold increase in effective batch size, and compute the InfoNCE loss over this enlarged set of embeddings. The classification results are reported in Table 3.
We also evaluate the performance of the asymmetric lower bound (18) (denoted ASYM) in place of the symmetrized objective (19). These results are also shown in Table 3.
From these experiments, we did not observe any significant differences when applying DistNCE (43), using the asymmetric loss, or sampling multiple embeddings per posterior. Based on these findings, we proceed with the basic VCL variants from the main text for all subsequent experiments.
Table 3: Classification accuracy on STL10 with different number of embedding generation from posterior. We report top-1 and top-5 accuracies of SimCLR, VSimCLR, SupCon, and VSupCon across the datasets with different $m$ and DistNCE (43).
Method STL10
Top1 Top5
SimCLR 60.44 95.80
VSimCLR ($m=1$) 60.11 92.00
VSimCLR ($m=4$) 57.86 88.29
VSimCLR ($m=16$) 59.13 92.85
VSimCLR ($m=64$) 56.91 86.63
VSimCLR with DistNCE (43) 36.54 80.25
VSimCLR (asym) 57.38 88.78
[…]
VSupCon ($m=1$) 75.76 96.99
VSupCon ($m=4$) 74.35 97.14
VSupCon ($m=16$) 76.11 98.39
VSupCon ($m=64$) 77.96 98.44